### v1

# Site-Decorated Model for Unconventional Frustrated Magnets with Ultranarrow Phase Crossover and Spin Reversal Transition

Abstract
The site-decorated Ising model is introduced to advance the understanding and physical realization of the recently discovered one-dimensional finite-temperature ultranarrow phase crossover in an external magnetic field, overcoming the complexity of the traditional bond-decorated models from geometric consideration. Furthermore, for higher-dimensional Ising models in the presence of an external magnetic field, while they remain unsolved, an exact solution about a novel spin-reversal transition—accessible by a slight change in temperature or the magnetic field, even in the weak field limit—is found to exist upon the site decoration. These results suggest a new route to energy-efficient applications in, e.g., data storage and processing, and call for materialization and device design with site decoration in, e.g., mixed $d$-$f$ compounds, optical lattices, or neural networks.
I Introduction
Finding new states with desirable physical properties and energy-efficient ways of phase transition to these states is a central problem in many research fields ranging from condensed matter physics and materials science to quantum information and mircroelectronics. For example, magnetization reversal in a ferromagnet or ferrimagnet is a basic operation in data storage and processing, but the traditional magnetic-field-driven spin flipping could suffer from considerable coercive force in those materials [1]; the current-induced spin–orbit torque has emerged as a candidate for more effective approaches to magnetization manipulation [2]. On the other hand, new states with fast response abound in frustrated magnets as spin ice, spin glass, spin liquid, spin vortex and skyrmion, etc., [3, 4, 5]. To date, the conventional frustration, termed as geometrical frustration, describes a macroscopic ground-state degeneracy in the spin arrangement caused by competing spin-spin interactions  [6]. For example, to geometrically frustrate the interaction between two spins, couple them to one additional spin [Figs. 1(a)(c)]: Six spin configurations are degenerate when the three antiferromagnetic bonds of the resulting spin triangle are equally strong. This approach is referred to as bond decoration (i.e., decorating the bond between the first two spins). Bond decoration with strong geometric frustration has been widely used to explore advanced systems with novel phase changes [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], such as the recently discovered finite-temperature ultranarrow phase crossover (UNPC) in one-dimensional Ising model, where a finite-temperature phase transition is forbidden but now found approachable by UNPC [11, 12, 13].
The purpose of this paper is to introduce a site-decorated model for UNPC and phase transition, where the conventional geometric frustration is apparently absent [Figs. 1(b)(d)]. The idea was stimulated by the UNPC phenomenon driven by a hidden “half ice, half fire” (HIHF) state in a one-dimensional (1D) bond-decorated ferrimagnet [13, 14]. This exotic state has a macroscopic degeneracy, but as the phase boundary between excited states, it is always hidden in the ground-state phase diagram. The concept of atomic-scale mixed ice-fire states offers a new materials-design guide [15]. Although the ferrimagnet studied was still a bond-decorated system, the spin-spin interactions are not competing at all when $J$ is ferromagnetic [Figs. 1(a)(c)]. Such irrelevance of geometric frustration induced ground-state degeneracy suggests that it be likely and timely to bypass the complexity of bond decoration in search for unconventional frustrated magnets that are much easier to be synthesized and understood.
Figure 1: Schematics of minimally decorated Ising models in one and two dimensions: (a)(c) bond decoration and (b)(d) site decoration. Green balls depict the original, type-a spins coupled by the interaction $J$ (red and orange bonds); gray balls depicts the decorated, type-b spins coupled to the backbone by the antiferromagnetic interaction $J_{b}<0$ (gray bonds). The spins’ magnetic moments are $\mu_{a}$ and $\mu_{b}$, respectively. The conventional geometric frustration is present [7, 8] and absent [13, 14] in the bond-decorated models (a)(c) for antiferromagnetic and ferromagnetic $J$, respectively. Such geometric frustration is clearly absent in the site-decorated models (b)(d).
Here, using the Ising model, I show exactly that the site-decorated ferrimagnet displays a similar HIHF-driven UNPC at fixed finite temperature in one dimension. Furthermore, for the two- and three-dimensional (2D and 3D) site-decorated Ising models in an external magnetic field, while they remain unsolved, I am able to rigorously reveal a HIHF-driven spin reversal transition. The spin reversal is accessible by a slight change in temperature or the magnetic field, even in the weak field limit, suggesting a new route to energy-efficient applications in data storage and processing for example. Implications on material and device design with, e.g., mixed $d$-$f$ compounds [1, 15], are addressed. These results establish site decoration as a new, feasible avenue to the exploration of novel physical properties in unconventional frustrated magnets.
II Results
II.1 One dimension
The minimally site-decorated Ising model in one dimension is shown in Fig. 1(b) and described by $H=H_{a}+H_{b}$, where
$H_{a}$ $=$ $-J\sum_{i=1}^{N}\sigma_{i}\sigma_{i+1}-h\mu_{a}\sum_{i=1}^{N}\sigma_{i},$ (1)
$H_{b}$ $=$ $-2J_{b}\sum_{i=1}^{N}\sigma_{i}b_{i}-h\mu_{b}\sum_{i=1}^{N}b_{i}.$ (2)
$H_{a}$ describes the backbone of the single chain with $\sigma_{i}=\pm 1$ (green balls, referred to as type-a spins) and $J>0$ the ferromagnetic interaction (red bonds). This is the model originally studied by Ising one century ago [16]. $H_{b}$ describes the decorated parts, where $b_{i}=\pm 1$ (gray balls, referred to as type-b spins) is coupled to the $i$th type-a spin with the antiferromagnetic interaction $02J_{b}<$ 0 (gray bonds). $h$ depicts the magnetic field, $\mu_{a}$ and $\mu_{b}$ the magnetic moments of type-a and type-b spins, respectively. The relationship of $\mu_{b}>\mu_{a}>0$ is used to represent ferrimagnetism—the model is invariant under the sign changes that satisfy $\mu_{a}\mu_{b}J_{b}<0$, e.g., $\mu_{a}=2$ and $\mu_{b}=-3$ in Sr3CuIrO6 [17, 18]. $N$ is the total number of the unit cell with $\sigma_{N+1}\equiv\sigma_{1}$, $b_{N+1}\equiv b_{1}$, viz., the periodic boundary condition.
In the calculation of the partition function $Z=\mathrm{Tr}\,e^{-\beta H}$ and the free energy per unit cell $f=-\frac{1}{N\beta}\ln Z$ in the thermodynamic limit $N\to\infty$, where $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the absolute temperature and $k_{\mathrm{B}}$ the Boltzmann constant, the type-b spins can be summed out exactly, resulting in $Z=e^{NA}\mathrm{Tr}\,e^{-\beta H_{\mathrm{eff}}}$ where
$A=\frac{1}{2}\left(\ln a+\ln b\right)$ (3)
with $a=2\cosh(\beta h\mu_{b}+2\beta J_{b})$ and $b=2\cosh(\beta h\mu_{b}-2\beta J_{b})$. $H_{\mathrm{eff}}$ is the effective Hamiltonian for the type-a spins:
$H_{\mathrm{eff}}=-J\sum_{i=1}^{N}\sigma_{i}\sigma_{i+1}-h_{\mathrm{eff}}\mu_{a}\sum_{i=1}^{N}\sigma_{i},$ (4)
which is of the same form as the undecorated Ising model defined in Equation (1)—with $h$ being replaced by a temperature-dependent effective magnetic field
$h_{\mathrm{eff}}=h+\frac{1}{2\beta\mu_{a}}\left(\ln a-\ln b\right).$ (5)
$h_{\mathrm{eff}}$ is independent of $J$. Compared with the bond-decorated Ising model, $J$ is not normalized in the site-decorated model, which greatly simplifies the analysis.
Figure 2: Density plot of (a) $h_{\mathrm{eff}}$ and (b) $\langle\sigma_{i}\rangle$ in the $h$-$T$ plane for the 1D site-decorated Ising model. The black solid line in (a) is the $h_{\mathrm{eff}}=0$ contour line, i.e., the $T_{0}$ line defined in Equation (7), which is the $\langle\sigma_{i}\rangle=0$ contour line in (b) as the phase boundary between the red and purple regions with $\langle\sigma_{i}\rangle=\pm 1$, respectively. As shown in (b), at zero temperature, there are two phases separated at $h=h_{c}\equiv 2(-J_{b})/\mu_{a}$: (i) the ferrimagnetic state with all type-a spins down and all type-b spins up for $h<h_{c}$, and (ii) the ferromagnetic state with all spins up for $h>h_{c}$. The dashed line at $h=h_{f}\equiv 2(-J_{b})/\mu_{b}$, which is the phase boundary of two excited phases there (the ferromagnetic state and the opposite ferrimagnetic state with all type-a spins up and all type-b spins down), indicates where the half-ice, half-fire state (where all type-a spins are ordered up and all type-b spins are disordered) hides. Here $J_{b}=-1$, $\mu_{a}=1$, and $\mu_{b}=4/3$.
The order parameter that allows for fast and accurate identification of UNPC is the magnetization of the type-a spins given by $\langle\sigma_{i}\rangle=-\frac{\partial f}{h\,\partial\mu_{a}}=-\frac{\partial(f+A/\beta)}{\partial(h_{\mathrm{eff}}\mu_{a})}$ [13, 14]:
$\langle\sigma_{i}\rangle=\frac{\sinh(\beta h_{\mathrm{eff}}\mu_{a})}{\sqrt{\sinh^{2}(\beta h_{\mathrm{eff}}\mu_{a})+e^{-4\beta J}}},$ (6)
which switches sign at the crossover temperature $T_{0}$ with the crossover width $2\delta T=2\left(\frac{\partial\langle\sigma_{i}\rangle}{\partial T}\right)^{-1}_{T=T_{0}}\propto e^{-2J/k_{\mathrm{B}}T_{0}}$. Apparently, $T_{0}$ can be determined by $h_{\mathrm{eff}}=0$ [10, 13, 14], yielding
$e^{4\beta J_{b}}=\frac{\sinh[\beta h(\mu_{b}-\mu_{a})]}{\sinh[\beta h(\mu_{b}+\mu_{a})]}.$ (7)
It has a finite-temperature solution for $\mu_{b}>\mu_{a}$, $J_{b}<0$, and $0<h<2(-J_{b})/\mu_{a}$. As shown in Fig. 2(a) for $h_{\mathrm{eff}}$, Fig. 2(b) for $\langle\sigma_{i}\rangle$, and End Matter for $\langle b_{i}\rangle$, the essential physics of UNPC driven by the hidden half-ice (type-a spins are fully ordered), half-fire (type-b spins are fully disordered) state in the bond-decorated Ising model [14] has been reproduced in the site-decorated model, where $\langle\sigma_{i}\rangle$ flips from $+1$ to $-1$ with $T_{0}$ fixed and $2\delta T$ narrowed exponentially as $J$ increases. As $h$ increases, both $T_{0}$ and $\delta T$ decrease until $T_{0}=0$ and $02\delta T=$ 0 at $h=2(-J_{b})/\mu_{a}$. The largest $T_{0}$ as a function of $h$ takes place at the weak field limit $h\to 0$,
$k_{\mathrm{B}}T^{\mathrm{max}}_{0}=4J_{b}\left(\ln\frac{1-\mu_{a}/\mu_{b}}{1+\mu_{a}/\mu_{b}}\right)^{-1}.$ (8)
Note that with the absolute zero field $h=0$, there is no sign change in $h_{\mathrm{eff}}$ for any finite temperature because $h_{\mathrm{eff}}\equiv 0$. This is consistent with the theorem that spontaneous UNPC does not exist in single-chain Ising models due to the spin-reversal symmetry associated with $h=0$ [11]. The minimal model for spontaneous UNPC is a decorated two-leg ladder Ising model [11, 12]. This abrupt change in the system’s behavior for $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
Figure 3: $T_{0}^{\mathrm{max}}$ (i.e., $T_{0}$ in the limit $h\to 0$) as a function of $\mu_{a}/\mu_{b}$. Here $J_{b}=-1$ and $\mu_{a}=1$. $T_{0}$ is independent of $J$.
As shown in Fig. 4, $T^{\mathrm{max}}_{0}$ quickly increases as $\mu_{a}/\mu_{b}$ decreases, a feature desirable for low-field, high-temperature applications; however, the crossover width also increases as $2\delta T\propto e^{-2J/k_{\mathrm{B}}T_{0}}/h$ in the weak $h$ limit. The divergence of $2\delta T$ at $h=0$ for fixed $J$ agrees with the aforementioned theorem, which however can be overcome in higher dimensions, as proven below.
Figure 4: The reversal of $\langle\sigma_{i}\rangle$ at $T_{0}$ as $h_{\mathrm{eff}}\to 0^{\pm}$ in the site-decorated square-lattice Ising model for $J=1$ with $T_{0}^{\mathrm{max}}<T_{c}^{*}\approx 2.269/k_{\mathrm{B}}$ (blue solid and dashed lines) and $J=0.88$ with $T_{0}^{\mathrm{max}}>T_{c}^{*}\approx 2/k_{\mathrm{B}}$ (red solid and dashed lines). Here $J_{b}=-1$, $\mu_{a}=1$, $\mu_{b}=4/3$ yielding $T_{0}^{\mathrm{max}}\approx 2.056/k_{\mathrm{B}}$.
II.2 Two dimension
Although the Ising model in two dimension with nonzero magnetic field remains unresolved, it is appreciated that site decoration affects only the effective magnetic field in the effective Hamiltonian and $h_{\mathrm{eff}}=0$ at $T_{0}$. Equations (3)-(5), (7), and (8) can be directly applied to the 2D cases. Therefore, one can reach the following exact result as a function of $h$.
For a site-decorated square lattice [Fig. 1(d)], at $T_{0}$ it reduces to a simple square lattice, which would undergo a paramagnetic-ferromagnetic phase transition at [19, 20]
$k_{\mathrm{B}}T_{c}^{*}=2J/\ln{(1+\sqrt{2})}\approx 2.2692J$ (9)
with the order parameter for $h_{\mathrm{eff}}\to 0$ given by [21, 22]
$\langle\sigma_{i}\rangle=\mathrm{sgn}(h_{\mathrm{eff}})\left[1-(\sinh 2\beta_{0}J)^{-4}\right]^{1/8}$ (10)
for $T_{0}<T_{c}^{*}$ and zero for $T_{0}>T_{c}^{*}$, where $\beta_{0}=1/(k_{\mathrm{B}}T_{0})$. Hence, the sign change of $h_{\mathrm{eff}}$ at $T_{0}$ results in a spin reversal transition. Recall that the sign change of $h_{\mathrm{eff}}$ can be achieved by a slight change in temperature or the magnetic field at the $T_{0}$ curve [the black line in Fig. 2(a)]. As shown in Fig. 4, there are two cases: (i) For $T_{0}^{\mathrm{max}}<T_{c}^{*}$ (blue solid and dashed lines), the spin reversal occurs in the whole region of $0<h<2(-J_{b})/\mu_{a}$. (ii) For $T_{0}^{\mathrm{max}}>T_{c}^{*}$ (red solid and dashed lines), the spin reversal has a threshold in $h$, whose corresponding $T_{0}$ equals $T_{c}^{*}$.
Note that $T_{c}^{*}$ is generally not the genuine Curie temperature $T_{c}$ of the system, which remains unresolved. $T_{c}^{*}$ is a value implied by $h_{\mathrm{eff}}=0$ and useful for determining which of the above two cases applies.
II.3 Three dimension
Likewise, since Equations (3)-(5), (7), and (8) can be directly applied to the site decoration in any dimension, the spin reversal transition as $h_{\mathrm{eff}}\to 0^{\pm}$ takes place in the site-decorated 3D Ising model in an external magnetic field as well with two cases similar to Fig. 4, although the exact expression of the spontaneous magnetization, like Equation (10) for a square lattice, is not available for the 3D Ising model.
III Discussion
Figure 5: $M$-side site decoration with (a) $M=2$ and (b) $M=4$, which effectively reduces the ratio of $\mu_{a}/\mu_{b}$ by a factor of $M$.
The present results have the following immediate impacts and stimulate a series of subsequent work:
I have not solved the critical outstanding problem of the 2D and 3D Ising models in an external magnetic field. I used the salient feature of $h_{\mathrm{eff}}=0$ in the site-decorated Ising models to exactly reveal the spin reversal transition at $T_{0}$ in the 2D and 3D Ising model in an external magnetic field. Furthermore, reexamining the 2D or 3D bond-decorated ferrimagnetic Ising models with ferromagnetic $J$ [Fig. 1(c)], the spin reversal transition occurs, too (not shown), though not as obvious. The present result calls for Monte Carlo simulation and micromagnetic calculation of the complete $h-T$ phase diagram of such decorated 2D and 3D Ising models.
Traditionally, spin reversal in a ferromagnet or ferrimagnet is achieved by applying a magnetic field in the opposite direction. This operation, e.g., in magneto recording, has to overcome the coercive force of the magnetic material. It is often needed to increase the local temperature to so close to the Curie temperature $T_{c}$, typically by laser illumination, that a small applied magnetic field can induce the magnetization reversal [1]. Therefore, the present spin reversal mechanism is attractive as it is accessible by a slight change in temperature or the magnetic field. The transition is driven by the hidden half-ice, half-fire state, in which the sublattice of the decorated, type-b spins provides a macroscopic degeneracy of $2^{N}$, causing the extremely sensitive, energy-efficient response of the type-a spins at $T_{0}$. It is worth mentioning that such response can occur for $h\to 0$, especially in two and three dimensions. The abrupt change in the system’s behavior for $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
For the materialization of the site-decorated models, mixed $d$-$f$ compounds are material candidates [1, 15], in which the type-a and type-b spins can be $3d$-$5d$ and $4f$ elements, respectively, as the wave functions of the $3d$-$5d$ electronic orbitals are more likely to overlap than the $4f$ orbitals. The $4f$ magnetic moments normally do not talk with each other and act as the decorated spins. Besides, it is not uncommon to find larger magnetic moments in the $4f$ elements than the $3d$-$5d$ ones to satisfy $\mu_{b}>\mu_{a}$. It is also natural to explore optical lattices [23] and neural networks [24, 25], among others, for they have been built to study other Ising models.
Another way to meet the criterion of $\eta\equiv\mu_{a}/\mu_{b}<1$ is to do multi-side ($M$-side) site decoration. $M=1$ in Figs. 1(b)(d). The examples of $M=2$ and $4$ are illustrated in Fig. 5. Then, the effective magnetic field becomes
$h_{\mathrm{eff}}=h+\frac{M}{2\beta\mu_{a}}\left(\ln a-\ln b\right).$ (11)
This results in the replacement of $\mu_{a}$ by $\mu_{a}/M$ in Equations (7) and (8), i.e., $\eta$ is effectively reduced by a factor of $M$. Consequently, $T_{0}$ could be increased dramatically.
Last but not least, the fact that site and bond decorations are geometrically compatible would significantly enlarge the possibility of functional materials design. The advancement in our knowledge of the complexity and vast possibility of the Ising model, a basic model for cooperative phenomena in various physical, biological, economical, and social systems, by the way, provides a fundamental playground for machine learning and AI for sciences.
Acknowledgements.
Brookhaven National Laboratory was supported by U.S. Department of Energy (DOE) Office of Basic Energy Sciences (BES) Division of Materials Sciences and Engineering under contract No. DE-SC0012704.