### v2

# Site-Decorated Model for Unconventional Frustrated Magnets: Ultranarrow Phase Crossover and Spin Reversal Transition

Abstract
The site-decorated Ising model is introduced to advance the understanding and experimental realization of the recently discovered one-dimensional finite-temperature ultranarrow phase crossover in an external magnetic field, while mitigating the geometric complexities of traditional bond-decorated models. Furthermore, although higher-dimensional Ising models in an external field remain unsolved, an exact solution for a novel spin-reversal transition—driven by an exotic, hidden ‘half-ice, half-fire’ state induced by site decoration—is derived. This transition, triggered by a slight variation in temperature or magnetic field even in the weak-field limit, offers a promising route toward energy-efficient applications such as data storage and processing. The results establish site decoration as a compelling new avenue for materials and device design, particularly in systems such as mixed $d$-$f$ compounds, optical lattices, and neural networks.
I Introduction
Finding novel states with desirable physical properties and energy-efficient phase transitions is a central challenge in various research fields, including condensed matter physics, materials science, quantum information, and microelectronics [1]. For example, magnetization reversal in ferromagnets and ferrimagnets is crucial for data storage and processing, yet traditional field-driven spin flipping suffers from considerable coercive forces [2], while current-induced spin–orbit torque has emerged as a more effective alternative [3].
Frustrated magnets host a rich variety of unconventional states with rapid response dynamics, including spin ice, spin glass, spin liquid, spin vortex, and skyrmions [4, 5, 6]. Traditionally, frustration—termed geometric frustration—arises from competing spin-spin interactions, leading to macroscopic ground-state degeneracy [7]. A well-established approach to engineering such frustration is bond decoration, where an additional spin is introduced along a bond [Figs. 1(a)(c)] [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Notably, a finite-temperature ultranarrow phase crossover (UNPC) was recently discovered in one-dimensional (1D) bond-decorated Ising models with short-range interactions, both in the absence [13, 14] and presence [15] of an external magnetic field. While finite-temperature phase transitions have been theoretically forbidden in this fundamental 1D statistical mechanics system for a century [1, 16, 17], UNPC now enables them to be approached arbitrarily closely [13, 14, 15].
The purpose of this Letter is to introduce site decoration as an alternative paradigm for exotic phases, phase transitions, and UNPC, where additional spins are placed on lattice sites rather than bonds, thereby eliminating conventional geometric frustration [Figs. 1(b)(d)]. This idea was inspired by the UNPC phenomenon driven by a hidden half-ice, half-fire (HIHF) state [18] and a 1/3-ice, 2/3-fire state [15] in 1D bond-decorated ferrimagnets. These ice-fire states exhibit macroscopic degeneracy yet remain concealed in the ground-state phase diagram, emerging only as phase boundaries between excited states. Their existence suggests that atomic-scale mixed ice-fire states could serve as a guiding principle for materials design [19]. While these ferrimagnets were bond-decorated, their spin-spin interactions did not induce geometric frustration since $J$ was ferromagnetic [Figs. 1(a)(c)]. This key observation indicates that bypassing the complexities of bond decoration is both feasible and advantageous, paving the way for site-decorated frustrated magnets that are easier to synthesize and understand.
Figure 1: Schematics of minimally decorated Ising models in one and two dimensions: (a)(c) bond decoration and (b)(d) site decoration. Green balls depict the backbone, type-a spins coupled by the interaction $J$ (red and orange bonds); gray balls depicts the decorated, type-b spins coupled to the backbone by the antiferromagnetic interaction $J_{ab}$ (gray bonds). The spins’ magnetic moments are $\mu_{a}$ and $\mu_{b}$, respectively. The conventional geometric frustration is present [9, 10] and absent [15, 18] in the bond-decorated models (a)(c) for antiferromagnetic and ferromagnetic $J$, respectively. Such geometric frustration is clearly absent in the site-decorated models (b)(d).
Using the Ising model, I demonstrate exactly that the site-decorated ferrimagnet exhibits an HIHF-driven UNPC at a fixed finite temperature in one dimension. Furthermore, while the two- and three-dimensional (2D and 3D) site-decorated Ising models in an external magnetic field remain unsolved, I rigorously reveal a HIHF-driven spin reversal transition. This spin reversal can be induced by a slight change in temperature or magnetic field, even in the weak-field limit, suggesting a new pathway for energy-efficient applications in data storage and processing. Implications for material and device design, particularly in mixed $d$-$f$ compounds [2, 19], are also discussed. These findings establish site decoration as a novel and viable framework for exploring unique physical properties in unconventional frustrated magnets.
II Results
II.1 One dimension
The minimally site-decorated Ising model in one dimension is shown in Fig. 1(b) and described by $H=H_{a}+H_{b}$, where
$H_{a}$ $=$ $-J\sum_{i=1}^{N}\sigma_{i}\sigma_{i+1}-h\mu_{a}\sum_{i=1}^{N}\sigma_{i},$ (1)
$H_{b}$ $=$ $-J_{ab}\sum_{i=1}^{N}\sigma_{i}b_{i}-h\mu_{b}\sum_{i=1}^{N}b_{i}.$ (2)
$H_{a}$ describes the backbone of the single chain with $\sigma_{i}=\pm 1$ (green balls, referred to as type-a spins) and $J>0$ the ferromagnetic interaction (red bonds). This is the model originally studied by Ising one century ago [16]. $H_{b}$ describes the decorated parts, where $b_{i}=\pm 1$ (gray balls, referred to as type-b spins) is coupled to the $i$th type-a spin with the antiferromagnetic interaction $J_{ab}<0$ (gray bonds). $h$ depicts the magnetic field, $\mu_{a}$ and $\mu_{b}$ the magnetic moments of type-a and type-b spins, respectively. The relationship of $\mu_{b}>\mu_{a}>0$ is used to represent ferrimagnetism—the model is invariant under the sign changes that satisfy $\mu_{a}\mu_{b}J_{ab}<0$, e.g., $\mu_{a}=2$ and $\mu_{b}=-3$ in Sr3CuIrO6 [20, 21]. $N$ is the total number of the unit cell with $\sigma_{N+1}\equiv\sigma_{1}$, $b_{N+1}\equiv b_{1}$, viz., the periodic boundary condition.
In the calculation of the partition function $Z=\mathrm{Tr}\,e^{-\beta H}$ and the free energy per unit cell $f=-\frac{1}{N\beta}\ln Z$ in the thermodynamic limit $N\to\infty$, where $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the absolute temperature and $k_{\mathrm{B}}$ the Boltzmann constant, the type-b spins can be summed out exactly for any dimension, resulting in $Z=e^{NA}\mathrm{Tr}\,e^{-\beta H_{\mathrm{eff}}}$ where
$A=\frac{1}{2}\left(\ln\boxed{+}+\ln\boxed{-}\right)$ (3)
with the shorthand notation ±$\boxed{\pm}=2\cosh(\beta h\mu_{b}\pm\beta J_{ab})$. $H_{\mathrm{eff}}$ is the effective Hamiltonian for the type-a spins:
$H_{\mathrm{eff}}=-J\sum_{i=1}^{N}\sigma_{i}\sigma_{i+1}-h_{\mathrm{eff}}\mu_{a}\sum_{i=1}^{N}\sigma_{i},$ (4)
which is of the same form as the undecorated Ising model defined in Equation (1)—with $h$ being replaced by a temperature-dependent effective magnetic field
$h_{\mathrm{eff}}=h+\frac{1}{2\beta\mu_{a}}\left(\ln\boxed{+}-\ln\boxed{-}\right).$ (5)
$h_{\mathrm{eff}}$ is independent of $J$. Unlike in the bond-decorated Ising model, $J$ is not normalized in the site-decorated model, significantly simplifying the analysis.
Figure 2: Density plot of (a) $h_{\mathrm{eff}}$ and (b) $\langle\sigma_{i}\rangle$ in the $h$-$T$ plane for the 1D site-decorated Ising model. The black solid line in (a) is the $h_{\mathrm{eff}}=0$ contour line, i.e., the $T_{0}$ line defined in Equation (7), which is the $\langle\sigma_{i}\rangle=0$ contour line in (b) as the phase boundary between the red and purple regions with $\langle\sigma_{i}\rangle=\pm 1$, respectively. As shown in (b), at zero temperature, there are two phases separated at $h=h_{c}\equiv-J_{ab}/\mu_{a}$: (i) the ferrimagnetic state with all type-a spins down and all type-b spins up for $h<h_{c}$, and (ii) the ferromagnetic state with all spins up for $h>h_{c}$. The dashed line at $h=h_{f}\equiv-J_{ab}/\mu_{b}$, which is the phase boundary of two excited phases there (the ferromagnetic state and the opposite ferrimagnetic state with all type-a spins up and all type-b spins down), indicates where the half-ice, half-fire state (where all type-a spins are ordered up and all type-b spins are disordered) hides. Here $J_{ab}=-2$, $\mu_{a}=1$, and $\mu_{b}=4/3$.
Figure 3: $T_{0}^{\mathrm{max}}$ (i.e., $T_{0}$ in the limit $h\to 0$) as a function of $\mu_{a}/\mu_{b}$. Here $J_{ab}=-2$ and $\mu_{a}=1$. $T_{0}$ is independent of $J$.
The order parameter that allows for fast and accurate identification of UNPC is the magnetization of the type-a spins given by $\langle\sigma_{i}\rangle=-\frac{\partial f}{h\,\partial\mu_{a}}$ [15, 18]:
$\langle\sigma_{i}\rangle=\frac{\sinh(\beta h_{\mathrm{eff}}\mu_{a})}{\sqrt{\sinh^{2}(\beta h_{\mathrm{eff}}\mu_{a})+e^{-4\beta J}}},$ (6)
which switches sign at the crossover temperature $T_{0}$. Then, according to Equation (6), $T_{0}$ can be determined by setting $h_{\mathrm{eff}}=0$, yielding the critical equation:
$\tanh(-\beta J_{ab})=\frac{\tanh(\beta h\mu_{a})}{\tanh(\beta h\mu_{b})},$ (7)
$T_{0}$ elegantly links four parameters of the Hamiltonian, while the remaining parameter, $J$, exponentially narrows the crossover width $2\delta T=2\left(\frac{\partial\langle\sigma_{i}\rangle}{\partial T}\right)^{-1}_{T=T_{0}}\propto e^{-2J/k_{\mathrm{B}}T_{0}}$. Equation (7) has a finite-temperature solution for $|\mu_{a}|<|\mu_{b}|$, $\mu_{a}\mu_{b}J_{ab}<0$, and $0<h<h_{c}$, where $h_{c}\equiv|J_{ab}/\mu_{a}|$.
As shown in Fig. 2(a) for $h_{\mathrm{eff}}$, Fig. 2(b) for $\langle\sigma_{i}\rangle$, and End Matter for $\langle b_{i}\rangle$, the essential physics of UNPC—driven by the hidden half-ice (type-a spins are fully ordered), half-fire (type-b spins are fully disordered) state in the bond-decorated Ising model [18]—has been reproduced in the site-decorated model. In this framework, $\langle\sigma_{i}\rangle$ flips from $+1$ to $-1$ at a fixed $T_{0}$, while $2\delta T$ narrows exponentially as $J$ increases.
As $h$ increases, both $T_{0}$ and $\delta T$ decrease, reaching $T_{0}=0$ and $02\delta T=$ 0 at $h_{c}$. The maximum $T_{0}$ as a function of $h$ occurs at the weak field limit $h\to 0$,
$k_{\mathrm{B}}T^{\mathrm{max}}_{0}=2J_{ab}\left(\ln\frac{1-\mu_{a}/\mu_{b}}{1+\mu_{a}/\mu_{b}}\right)^{-1}.$ (8)
Note that with the absolute zero field ($h=0$), there is no sign change in $h_{\mathrm{eff}}$ because $h_{\mathrm{eff}}=0$ at any finite temperature. This is consistent with the theorem that spontaneous UNPC does not exist in single-chain Ising models due to the spin-reversal symmetry associated with $h=0$ [13]. The minimal model for spontaneous UNPC is a decorated two-leg ladder Ising model [13, 14]. This abrupt change in the system’s behavior between $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
As shown in Fig. 4, $T^{\mathrm{max}}_{0}$ rapidly increases as $\mu_{a}/\mu_{b}$ decreases, a feature desirable for low-field, high-temperature applications. However, the crossover width also increases as $2\delta T\propto e^{-2J/k_{\mathrm{B}}T_{0}}/h$ in the weak $h$ limit. The divergence of $2\delta T$ at $h=0$ for fixed $J$ agrees with the aforementioned theorem; however, this limitation can be overcome in higher dimensions, as proven below.
Figure 4: The reversal of $\langle\sigma_{i}\rangle$ at $T_{0}$ as $h_{\mathrm{eff}}\to 0^{\pm}$ in the site-decorated square-lattice Ising model for $J=1$ with $T_{0}^{\mathrm{max}}<T_{c}^{*}\approx 2.269/k_{\mathrm{B}}$ (blue solid and dashed lines) and $J=0.88$ with $T_{0}^{\mathrm{max}}>T_{c}^{*}\approx 2/k_{\mathrm{B}}$ (red solid and dashed lines). Here $J_{ab}=-2$, $\mu_{a}=1$, $\mu_{b}=4/3$ yielding $T_{0}^{\mathrm{max}}\approx 2.056/k_{\mathrm{B}}$.
II.2 Higher dimensions
Although 2D and 3D Ising models with a nonzero magnetic field remains unresolved, it has been proven for any dimension that site decoration yields $h_{\mathrm{eff}}=0$ at $T_{0}$ as a function of $h$ [the black line in Fig. 2(a)], while keeping $J$ unchanged. Equations (3)-(5), (7), and (8) are directly applicable to the 2D and 3D cases. Consequently, the following exact result can be derived as a function of $h$.
For a site-decorated square lattice [Fig. 1(d)], the system reduces to a simple square lattice with $h_{\mathrm{eff}}=0$ at $T_{0}$, which would undergo a spontaneous paramagnetic-ferromagnetic phase transition at [22, 23]
$k_{\mathrm{B}}T_{c}^{*}=2J/\ln{(1+\sqrt{2})}\approx 2.2692J$ (9)
with the order parameter for $h_{\mathrm{eff}}\to 0$ given by [24, 17]
$\langle\sigma_{i}\rangle=\mathrm{sgn}(h_{\mathrm{eff}})\left[1-(\sinh 2\beta_{0}J)^{-4}\right]^{1/8}$ (10)
for $T_{0}<T_{c}^{*}$ and zero for $T_{0}>T_{c}^{*}$, where $\beta_{0}=1/(k_{\mathrm{B}}T_{0})$.
Hence, the sign change of $h_{\mathrm{eff}}$ at $T_{0}$ induces a spin reversal transition. Recall that this sign change can be triggered by a slight variation in temperature or magnetic field along the $T_{0}$ curve [the black line in Fig. 2(a)]. As shown in Fig. 4, two cases arise: (i) If $T_{0}^{\mathrm{max}}<T_{c}^{*}$ (blue solid and dashed lines), spin reversal occurs throughout the entire region $0<h<h_{c}$. (ii) If $T_{0}^{\mathrm{max}}>T_{c}^{*}$ (red solid and dashed lines), spin reversal has a threshold in $h$, where the corresponding $T_{0}$ equals $T_{c}^{*}$.
Note that $T_{c}^{*}$ is generally not the true Curie temperature $T_{c}$ of the system, which remains unresolved. Instead, $T_{c}^{*}$ is a value derived for $h_{\mathrm{eff}}=0$ and serves as the criterion for determining which of the two cases applies.
Similarly, the spin reversal transition at $h_{\mathrm{eff}}=0$ ($h\neq 0$) also occurs in other site-decorated 2D and 3D Ising models under an external magnetic field. This transition follows two cases analogous to Fig. 4, with exact results calculable for the anisotropic rectangular, anisotropic triangular [25], honeycomb, and Kagomé [26] lattices, where the exact expression for spontaneous magnetization—akin to Equation (10) for a square lattice—is known.
III Discussion
Figure 5: $M$-side site decoration with (a) $M=2$ and (b) $M=4$, which effectively reduces the ratio of $\mu_{a}/\mu_{b}$ by a factor of $M$.
The present results have immediate implications and inspire a series of subsequent studies, as outlined below.
I have not yet solved the long-standing problem of the 2D and 3D Ising models in a magnetic field. However, by leveraging the key feature of $h_{\mathrm{eff}}=0$ in the site-decorated Ising models, I have exactly revealed the spin reversal transition at $T_{0}$ in the 2D and 3D Ising models under an external field. Furthermore, a reexamination of 2D and 3D bond-decorated models with ferromagnetic $J$ [Fig. 1(c)] also reveals a spin reversal transition (not shown), though it is less obvious. These findings highlight the need for Monte Carlo simulations [27] to map the complete $h-T$ phase diagram of such decorated 2D and 3D Ising models.
Traditionally, spin reversal in ferromagnets and ferrimagnets is induced by applying a magnetic field in the opposite direction. In applications such as magnetic recording, overcoming the coercive force often requires increasing the local temperature close to the Curie temperature $T_{c}$, typically through laser illumination, so that a small applied field can trigger magnetization reversal [2]. By contrast, the spin reversal mechanism presented here is particularly attractive, as it can be achieved through a slight change in temperature or magnetic field. This transition is driven by the hidden half-ice, half-fire state, in which the decorated type-b sublattice provides a macroscopic degeneracy of $2^{N}$, resulting in an extremely sensitive and energy-efficient response of the type-a spins at $T_{0}$. Notably, this response persists as $h\to 0$, particularly in two and three dimensions. The abrupt contrast in system behavior between $h=0$ and $h\to 0$ is rare and could serve as a method for measuring absolute zero magnetic field.
For the experimental realization of site-decorated models, mixed $d$-$f$ or $3d$-$5d$ compounds are promising candidates [2, 19]. In these systems, the type-a and type-b spins correspond to $d$ and $4f$ elements, respectively, since the electronic wavefunctions of $d$ orbitals exhibit greater overlap than those of $4f$ orbitals. The weak interaction between $4f$ moments allows them to act as decorated spins, and the typically larger magnetic moments of $4f$ elements compared to $d$ ones naturally satisfy the condition $\mu_{b}>\mu_{a}$. A similar argument applies to $3d$-$5d$ compounds. Additionally, optical lattices [28] and neural networks [29, 30], both of which have been employed to study other Ising models, present alternative platforms for exploring site-decorated systems.
Another way to satisfy the condition $\mu_{a}/\mu_{b}<1$ is through multi-side ($M$-side) site decoration. Figs. 1(b)(d) represent $M=1$. The examples of $M=2$ and $4$ are illustrated in Fig. 5. Then, the effective magnetic field becomes
$h_{\mathrm{eff}}=h+\frac{M}{2\beta\mu_{a}}\left(\ln\boxed{+}-\ln\boxed{-}\right).$ (11)
This results in the replacement of $\mu_{a}$ by $\mu_{a}/M$ in Equations (7) and (8), i.e., $\mu_{a}/\mu_{b}$ is effectively reduced by a factor of $M$. Consequently, $T_{0}$ can increase dramatically.
Finally, the geometric compatibility of site and bond decorations significantly expands the potential for functional materials design. Moreover, advancing our understanding of the complexity and vast possibilities of the Ising model—a fundamental framework for cooperative phenomena across physical, biological, economic, and social systems—provides a crucial foundation for machine learning and AI applications in science.
IV Summary
In summary, site decoration has been introduced as a novel scheme for frustrated magnets, enabling exotic phase-switching phenomena in any dimension and yielding a rare exact result for 2D and 3D systems in a magnetic field. These discoveries establish a new paradigm for designing materials and devices with extreme sensitivity, where a hidden frustrated state—the half-ice, half-fire state—drives an abrupt spin reversal even under an infinitesimal applied field.
Acknowledgements.
Brookhaven National Laboratory was supported by U.S. Department of Energy (DOE) Office of Basic Energy Sciences (BES) Division of Materials Sciences and Engineering under contract No. DE-SC0012704.